OpenStudy (anonymous):

Identify whether the series summation of 15 open parentheses 4 close parentheses to the i minus 1 power from 1 to infinity is a convergent or divergent geometric series and find the sum, if possible.

OpenStudy (anonymous):

\[\sum_{\infty}^{i=1}15(4)^i-1\]

OpenStudy (anonymous):

It's suppose to be 15(4)^i=1

OpenStudy (anonymous):

This is a convergent geometric series. The sum is –5. This is a divergent geometric series. The sum is –5. This is a convergent geometric series. The sum cannot be found. This is a divergent geometric series. The sum cannot be found.

OpenStudy (anonymous):

I got A :)

jimthompson5910 (jim_thompson5910):

the exponent is `i-1` ??

OpenStudy (anonymous):

Whether you meant \(5(4)^{i+1}\) or \(5(4)^{i-1}\), consider what happens as \(i\to\infty\). \[\lim_{i\to\infty}5(4)^{i}=\infty\neq0\] What do you know about series of the form \(\sum a_n\) for which \(\lim\limits_{n\to\infty}a_n\neq0\)?

OpenStudy (anonymous):

Yes the exponent is i-1, and I think it's divergent

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

And I couldnt find the sum so D!

jimthompson5910 (jim_thompson5910):

Yep whenever it's divergent, the infinite number of terms don't sum to one fixed number.